Please help to answer in correct units
STOICHIOMETRY Daysia Using molarity to find solute mass and solution volume A chemist adds 350.0 mL of a 3.1 mol/L calcium bromide (CaBr,) solution to a reaction flask. Calculate the mass in kilograms of calcium bromide the chemist has added to the flask. Round your answer to 2 significant digits. Explanation Check
Answer #1

Molarity = Mole/Volume (in L)

Mole = volume × Molarity (350mL= 0.35 L)

= 0.35×3.1

= 1.085 mol

Mass = mol × molar mass

= 1.085 × 200 = 217 g

Mass in Kg = 0.22 kg